OpenStudy (anonymous):

I need help on number 74

OpenStudy (yoloshroom):

Number 74?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Well what's the question for number 74 :p

OpenStudy (yoloshroom):

Yeah XD

OpenStudy (anonymous):

Sorry I guess the picture didn't get posted it is f(x)=-3x+1 Find the difference quotient of f;that is, find [F(x+h)-f(x)]/h,h cannot equal 0

OpenStudy (ybarrap):

$$ \href{ https:///en.wikipedia.org/wiki/74_(number)}{74} $$

OpenStudy (anonymous):

@yoloshroom

OpenStudy (ybarrap):

$$ f(x)=-3x+1\\ \frac{f(x+h)-f(x)}{h}\\ \frac{\left( -3(x+h)+1\right)-\left(-3x+1\right)}{h}\\ $$ Can you simplify this?

OpenStudy (anonymous):

Yes thank you